given f(x)=5x^2-3x+1, find f(-2)
A). 15
B). 27
C). -13
D). -25

Found 2 solutions by stanbon, Nate:
Answer by stanbon(75887)

f(x)=5x^2-3x+1, find f(-2)
f(-2)= 5(-2)^2-3(-2)+1
=5(4)+6 +1
=27
Cheers,
Stan H.

Answer by Nate(3500)

f(x) = 5x^2 - 3x + 1
f(-2) = 5(-2)^2 - 3(-2) + 1
f(-2) = 20 + 6 + 1
f(-2) = 27
(B)